OpenStudy (anonymous):

How many solutions does 3x+6 = -1 - 3 + 4x have ? A. 2 B. 0 C. 1 D. Infinitely many

OpenStudy (lynfran):

3x+6=-4+4x 6+4=4x-3x 10=x only 1 solution do u understand?

OpenStudy (anonymous):

That's what i thought i just didnt know if i was correct.

OpenStudy (lynfran):

ok